Circular No. 6536

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997C IN NGC 3160 W. Li, Q. Qiao, Y. Qiu, D. Lu, and J. Hu, Beijing Astronomical Obser-

vatory (BAO), on behalf of the BAO Supernova Survey, report the discovery of an apparent supernova (mag ∼ 17.5) on unfiltered CCD images taken on Jan. 13.74, 13.85, and 13.93 UT with the 0.60-m BAO reflector. SN 1997C is located at α = 10^{h}13^{m}56^{s}.18, δ = +38^{o}49′00′′.5 (equinox 2000.0), which is 17′′.7 east and 96′′.1 south of the nucleus of NGC 3160. Nothing is present on images of NGC 3160 taken on 1996 Dec. 18 (limiting mag ∼ 19.0). A spectrum (range 380–900 nm, resolution 40 nm/mm) obtained with the BAO 2.16-m telescope at Xinglong station shows broad emission and absorption lines typical of a type-Ic supernova, ∼ 3–4 weeks after maximum. The spectrum is very similar to that of SN 1987M, ∼ 29 days after maximum (Wheeler and Benetti 1995, preprint). Strong spectral lines include the Ca II infrared triplet and H and K lines; O I at 760 nm; Na I D at 585 nm; Si II at 636 nm; and Fe II at 424, 448, 490, and 502 nm. There is no distinct evidence of H Balmer lines. The expansion velocity of the supernova envelope is ∼ 12 300 km/s when measured from the Ca II triplet, 15 000 km/s when measured from O I 760-nm, and 15 900 km/s when measured from the Ca II H and K lines. CCD frames taken on Jan. 14 give an unfiltered magnitude of ∼ 17.4.

GX 1+4 R. B. Wilson, Marshall Space Flight Center, NASA; and D. Chakra-

barty, Massachusetts Institute of Technology, report for the Compton Observatory BATSE Team: “Pulsations from the accreting x-ray pulsar GX 1+4, undetected (15-mCrab upper limit) in the energy range 20–50 keV since 1996 Aug. 18 (IAUC 6478), are once again being observed. The source first reappeared in data summed over Nov. 29.0–Dec. 4.0 UT, at a barycentric pulse frequency of 8.00122 ± 0.00037 mHz. The pulsations have continued through 1997 Jan. 12.5, with a frequency on that date of 7.9736 ± 0.0005 mHz and a pulsed flux of 23 ± 3 mCrab (20–50 keV). The mean pulse frequency derivative since redetection is (−7.7 ± 0.1) × 10^{−12} Hz s^{−1}, about twice the spindown rate observed in 1996 May–June; such values have previously been observed from the source. The inferred mean spindown rate during the undetected interval from Aug. 18 to Dec. 1.5 is (−3.35 ± 0.07) × 10^{−12} Hz s^{−1}. Follow-up optical and infrared observations of the mass donor V2116 Oph are strongly encouraged.”

1997 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green